# Electronic structure, spin-orbit interaction and electron–phonon coupling of triangular adatom lattices on semiconductor substrates

### v1

@@ III Results II: Phononic structure
Ultimately we are interested in the EPC strength; in order to compute it we need to derive the phonon spectra first. For Sn/Si(111) that was previously done to explore the consequences for the correlated many-body phases Zahedifar and Kratzer (2019); Wolf et al. (2022). Here we consistently ignore many-body phases such as magnetic or charge order, which would drastically affect the phonon bandstructures Zahedifar and Kratzer (2019). We use the software package phonopy Togo et al. (2023); Togo (2023) along with VASP to compute the phonon spectra, as shown in Fig. 4 a-d. We emphasize that the phonon computations are performed without spin-orbit coupling; the spin-orbit case is considered further below. We have colored the phonon eigenvalues according to the adatom weight in the eigenmodes, so that the contribution of surface adatoms is emphasized. Apart from the differences of frequencies $\omega$ at the $K$ and $M$ points, all four phonon spectra look essentially the same up to some global energy scaling. In particular, one can clearly see that the three Sn acoustic phonon bands disperse inside a gap into the Si(111) phonon dispersion. One should note the differing energy scales for X/Si(111) and X/SiC(0001) dispersions, showing higher energy SiC bulk modes.
Figure 4: Phonon dispersions for a Sn/Si(111), b Pb/Si(111), c Sn/SiC(0001) and d Pb/SiC(0001). As with the electronic dispersions DFT+$U$ has been used for phonon calculations. The colour of each point represents the adatom contribution to the mode (see color bar). Effective band structures for e Sn/Si(111), f, g Pb/Si(111), h Sn/SiC(0001) and i Pb/SiC(0001) for electron-phonon coupling calculations with 0.2 Å displacement amplitude. The selected phonon mode is indicated by the arrows in panels a-d. Electron-phonon coupling calculations have a reduced Brillouin zone due to the supercell required to encompass the effects of a given phonon mode. The band structure in the reduced Brillouin zone is thus unfolded onto the Brillouin zone of the original unit cell as an effective band structure. The band structure of the unperturbed unit cell is plotted in red for reference. The size of each blue dot represents band unfolding weight.
@@ IV Results III: Electron-Phonon Coupling
For a given mode we first create a supercell large enough to realise the full size of the phonon. At the $M$ and $K$ modes this corresponds to $2\times 2$ and $3\times 3$ supercells containing 4 (160) and 9 (360) adatoms (atoms), respectively. The supercell atoms are then perturbed according to the chosen phonon mode eigenvector and the associated band structure is calculated. Due to the reduction in Brillouin zone size as we increase the size of our supercell, we must unfold the band structure onto the unit cell Brillouin zone to obtain an effective band structure. For each mode we perturb atomic positions by a maximum of $\Delta Q=0.1\rm{\AA}$ and $\Delta Q=0.2\rm{\AA}$. It is important to choose only small displacement amplitudes to stay within the linear response regime. From the effective band structures we then calculate the deformation potentials by comparing the gap opening $2\Delta E_{k}$ with $\Delta Q$. In Fig. 4 e-i we show unfolded effective band structures with $\Delta Q=0.2\rm{\AA}$ for different phonon modes and materials. For band structures coupled to modes with both phonon wavevector $q=M$ and $q=K$ we find distinct gap openings along the $\Gamma K$ and $M\Gamma$ paths. For each material and coupled mode we calculate deformation potentials for the different gap openings, the largest of these being used to calculate $\lambda$. Due to the large number of atoms required in the EPC calculations it was unfeasible for us to perform all these calculations including relativistic effects.
Material Mode $\lambda$ $\omega\,(meV)$ $\mathcal{D}\,(eV/\text{\AA})$ $N(E_{F})$
@@ VI Discussion
Tab. 5 clearly rules out strong EPC as known e.g.,  from MgB2An and Pickett (2001). The X/SiC(0001) materials have such a small $\lambda$ that any phonon-mediated superconductivity is out of the picture. Only Sn/Si(111) with its phonon mode at the $M$ point has a somewhat larger EPC strength $\lambda=0.135$. To further analyze if such a value of $\lambda$ could possibly be responsible for the observed superconducting transition at $T_{c}=4.7$K, we try to estimate $T_{c}$ via the Allen-Dynes equationAllen and Dynes (1975) $T_{c}=\langle\omega\rangle/1.20\exp{(-\frac{1.04(1+\lambda)}{\lambda-\mu^{\ast}(1+0.62\lambda)})}$. Assuming $\omega\approx 100$K and a Coulomb pseudopotential in the typical range $\mu^{\ast}\sim 0.05-0.2$ Carbotte (1990), we find for Sn/Si(111) $T_{c}\sim 10^{-15}$ $\approx 0$. It is worth emphasizing that the deformation potential method used here to estimate $\lambda$ replaces the momentum-dependence of EPC with its value at the respective high symmetry point, which is likely to overestimate EPC and thus overestimate $T_{c}$. Moreover, if we follow the ideas of Ref. Koch et al., 1999 to quantify $\mu^{\ast}$ we would end up with an even larger value than $\mu^{\ast}\sim 0.05-0.2$, causing a further reduction of $T_{c}$. Despite the approximate character of estimating $T_{c}$ employed in this paper, we conclude that it does not appear to be plausible that EPC could be responsible for the observed superconductivity in Sn/Si(111) Ming et al. (2018, 2023).
@@ Acknowledgements.
We acknowledge discussions with Steve Johnston and Hanno Weitering. S.R. acknowledges support from the Australian Research Council through Grant No. DP200101118 and DP240100168. This research was undertaken using resources from the National Computational Infrastructure (NCI Australia), an NCRIS enabled capability supported by the Australian Government.
@@ Appendix A Electronic structure
Figure 6: Schematic of the nearest-neighbors included in the surface state Wannierisation. Symmetry planes preserved by the substrate are also shown (cf. Fig. 2).
@@ A.1 Group theory decomposition of Hamiltonian
Sn/Si(111)
F. Adler 2019Adler et al. (2019) (LDA) $52.2$ $-0.3889$ $0.1418$ $-0.0268$
Li 2013Li et al. (2013) (LDA) $52.7$ $-0.3881$ $0.1555$ $-0.0228$ $-0.0318$
D. I. BadrtdinovBadrtdinov et al. (2016) 2016 (PBE) $43.51$ $-0.4365$ $0.1271$ $0.0198$ $0.0032$
P. Hansmann 2013Hansmann et al. (2013) (GW+DMFT) $42.0$ $-0.4762$ $0.2381$
Pb/Si(111)
F. Adler 2019 (LDA) $58.5$ $-0.3829$ $0.1248$ $-0.0239$
D. I. Badrtdinov 2016 (PBE) $41.32$ $-0.4635$ $0.4037$ $0.0506$ $0.0027$
P. Hansmann 2013 (GW+DMFT) $42.0$ $-0.4762$ $0.2381$
Sn/SiC(0001)
D. I. BadrtdinovBadrtdinov et al. (2018) 2018 (PBE) $27.94$ $-0.5140$ $0.2645$ $0.0537$ $0.0043$
S. GlassGlass et al. (2015) 2015 (LDA) $27.3$
Pb/SiC(0001)
@@
where ${\Gamma_{ij}=-\rm{sgn}(\sin(6\theta_{ij}))}$. When the symmetry is reduced to $C_{3v}$, the valley-alternating versions the $C_{6v}$ symmetric terms are allowed. The forms of these were already given in Eq. 14-16, and can be extended to the radial Rashba term Eq. 17 by an analagous inclusion of the factor $\nu_{ij}$.
@@ A.2 Full band structure
In Fig. 7 we show the same band structure as in Fig. 1 in the main text but over a wider energy range. DFT+$U$ has been used to increase the semiconductor band gaps to match experiment, which is reflected in the figure. One can see that the surface bands disperse through the semiconductor band gap, and we note that there is no energy overlap between the bulk bands and the surface bands in any of the materials.
@@ Appendix B Electron-phonon coupling
In Tab. 7 we compare EPC calculations performed with and without DFT+$U$ for the X/Si(111) materials, in-plane phonon modes, and the effect of SOC in phonon calculations. We find that the inclusion of DFT+$U$ causes a reduction in the EPC strength $\lambda$. Notably, the deformation potential $\mathcal{D}$ is mostly unaffected, with the largest change $\Delta\mathcal{D}=0.010\,$eV/Å  in Pb/Si(111) coupled to the band 1 $K$ mode. The source of this reduction can be mostly attributed to the general increase in phonon mode frequency seen in both materials for a given mode, as $\lambda\sim 1/\omega^{2}$. For both Sn/Si(111) and Sn/SiC(0001) we test phonon modes with eigenvector corresponding to in-plane motion. For these calculations we find no discernible change in the effective band structure as compared to the unperturbed unit cell and thus minimal coupling. It is unsurprising that out-of-plane motion will have the largest effect on the band structures when considering hybridisation of Wannier functions with the bulk. As discussed in the main text, the reduction in hybridisation corresponds to significant changes in tight-binding hopping parameters. Changes in the adatom height above the substrate should have the largest effect on hybridisation, while in-plane motion should leave this largely unchanged.
Material Mode $\lambda$ $\omega\,(meV)$ $\mathcal{D}\,(eV/\text{\AA})$

### v2

@@ III Results II: Phononic structure
Ultimately we are interested in the EPC strength; in order to compute it we need to derive the phonon spectra first. For Sn/Si(111) that was previously done to explore the consequences for the correlated many-body phases Zahedifar and Kratzer (2019); Wolf et al. (2022). Here we consistently ignore many-body phases such as magnetic or charge order, which would drastically affect the phonon bandstructures Zahedifar and Kratzer (2019). We use the software package phonopy Togo et al. (2023); Togo (2023) along with VASP to compute the phonon spectra, as shown in Fig. 4 a-d. We emphasize that the phonon computations are performed without spin-orbit coupling; the spin-orbit case is considered further below. We have colored the phonon eigenvalues according to the adatom weight in the eigenmodes, so that the contribution of surface adatoms is emphasized. Apart from the differences of frequencies $\omega$ at the $K$ and $M$ points, all four phonon spectra look essentially the same up to some global energy scaling. In particular, one can clearly see that the three Sn acoustic phonon bands disperse inside a gap into the Si(111) phonon dispersion. One should note the differing energy scales for X/Si(111) and X/SiC(0001) dispersions, showing higher energy SiC bulk modes due to the lower average atomic mass in SiC compared to elemental Si.
Figure 4: Phonon dispersions for a Sn/Si(111), b Pb/Si(111), c Sn/SiC(0001) and d Pb/SiC(0001). As with the electronic dispersions DFT+$U$ has been used for phonon calculations. The color of each point represents the adatom contribution to the mode (see color bar). Effective band structures for e Sn/Si(111), f, g Pb/Si(111), h Sn/SiC(0001) and i Pb/SiC(0001) for electron-phonon coupling calculations with 0.2 Å displacement amplitude. The selected phonon mode is indicated by the arrows in panels a-d. Electron-phonon coupling calculations have a reduced Brillouin zone due to the supercell required to encompass the effects of a given phonon mode. The band structure in the reduced Brillouin zone is thus unfolded onto the Brillouin zone of the original unit cell as an effective band structure. The band structure of the unperturbed unit cell is plotted in red for reference. The size of each blue dot represents band unfolding weight.
@@ IV Results III: Electron-Phonon Coupling
For a given mode we first create a supercell large enough to realize the full size of the phonon. At the $M$ and $K$ modes this corresponds to $2\times 2$ and $3\times 3$ supercells containing 4 (160) and 9 (360) adatoms (atoms), respectively. The supercell atoms are then perturbed according to the chosen phonon mode eigenvector and the associated band structure is calculated. Due to the reduction in Brillouin zone size as we increase the size of our supercell, we must unfold the band structure onto the unit cell Brillouin zone to obtain an effective band structure. For each mode we perturb atomic positions by a maximum of $\Delta Q=0.1\rm{\AA}$ and $\Delta Q=0.2\rm{\AA}$. It is important to choose only small displacement amplitudes to stay within the linear response regime. From the effective band structures we then calculate the deformation potentials by comparing the gap opening $2\Delta E_{k}$ with $\Delta Q$. In Fig. 4 e-i we show unfolded effective band structures with $\Delta Q=0.2\rm{\AA}$ for different phonon modes and materials. For band structures coupled to modes with both phonon wavevector $q=M$ and $q=K$ we find distinct gap openings along the $\Gamma K$ and $M\Gamma$ paths. For each material and coupled mode we calculate deformation potentials for the different gap openings, the largest of these being used to calculate $\lambda$. Due to the large number of atoms required in the EPC calculations it was unfeasible for us to perform all these calculations including relativistic effects.
Material Mode $\lambda$ $\omega\,{\rm(meV)}$ $\mathcal{D}\,({\rm eV}/\text{\AA})$ $N(E_{F})$
@@ VI Discussion
Tab. 5 clearly rules out strong EPC as known e.g.,  from MgB2An and Pickett (2001). The X/SiC(0001) materials have such a small $\lambda$ that any phonon-mediated superconductivity is out of the picture. Only Sn/Si(111) with its phonon mode at the $M$ point has a somewhat larger EPC strength $\lambda=0.135$. To further analyze if such a value of $\lambda$ could possibly be responsible for the observed superconducting transition at $T_{c}=4.7$K, we try to estimate $T_{c}$ via the Allen-Dynes equationAllen and Dynes (1975) $T_{c}=\langle\omega\rangle/1.20\exp{(-\frac{1.04(1+\lambda)}{\lambda-\mu^{\ast}(1+0.62\lambda)})}$. Assuming $\omega\approx 100$K and a Coulomb pseudopotential in the typical range $\mu^{\ast}\sim 0.05-0.2$ Carbotte (1990), we find for Sn/Si(111) $T_{c}\sim 10^{-15}$ $\approx 0$.
It is worth emphasizing that the deformation potential method used here to estimate $\lambda$ replaces the momentum-dependence of EPC with its value at the respective high symmetry point, which is likely to overestimate EPC and thus overestimate $T_{c}$. Moreover, if we follow the ideas of Ref. Koch et al., 1999 to quantify $\mu^{\ast}$ we would end up with an even larger value than $\mu^{\ast}\sim 0.05-0.2$, causing a further reduction of $T_{c}$. Despite the approximate character of estimating $T_{c}$ employed in this paper, we conclude that it does not appear to be plausible that EPC could be responsible for the observed superconductivity in Sn/Si(111) Ming et al. (2018, 2023).
@@ VII Outlook
Data availability
All data needed to evaluate the conclusions in the paper are present in the paper and are available for download at Ref. Marchetti et al., 2025.
@@ Acknowledgements.
We acknowledge discussions with Steve Johnston, Hanno Weitering and Hua Chen. S.R. acknowledges support from the Australian Research Council through Grant No. DP200101118 and DP240100168. This research was undertaken using resources from the National Computational Infrastructure (NCI Australia), an NCRIS enabled capability supported by the Australian Government.
@@ Appendix A Electronic structure
Figure 6: Schematic of the nearest-neighbors included in the surface state Wannierization. Symmetry planes preserved by the substrate are also shown (cf. Fig. 2).
@@ A.1 Group theory decomposition of Hamiltonian
where ${\Gamma_{ij}=-\rm{sgn}(\sin(6\theta_{ij}))}$. When the symmetry is reduced to $C_{3v}$, the valley-alternating versions of the $C_{6v}$ symmetric terms are allowed. The forms of these were already given in Eq. 14-16, and can be extended to the radial Rashba term Eq. 17 by an analogous inclusion of the factor $\nu_{ij}$.
@@
Sn/Si(111)
F. Adler 2019Adler et al. (2019) (LDA) $52.2$ $-0.3889$ $0.1418$ $-0.0268$
Li 2013Li et al. (2013) (LDA) $52.7$ $-0.3881$ $0.1555$ $-0.0228$ $-0.0318$
D. I. Badrtdinov 2016Badrtdinov et al. (2016) (PBE) $43.51$ $-0.4365$ $0.1271$ $0.0198$ $0.0032$
P. Hansmann 2013Hansmann et al. (2013) (GW+DMFT) $42.0$ $-0.4762$ $0.2381$
Pb/Si(111)
F. Adler 2019 (LDA) $58.5$ $-0.3829$ $0.1248$ $-0.0239$
D. I. Badrtdinov 2016 (PBE) $41.32$ $-0.4635$ $0.4037$ $0.0506$ $0.0027$
P. Hansmann 2013 (GW+DMFT) $42.0$ $-0.4762$ $0.2381$
Sn/SiC(0001)
D. I. Badrtdinov 2018Badrtdinov et al. (2018) (PBE) $27.94$ $-0.5140$ $0.2645$ $0.0537$ $0.0043$
S. Glass 2015Glass et al. (2015) (LDA) $27.3$
Pb/SiC(0001)
@@ A.2 Full band structure
In Fig. 7 we show the same band structure as in Fig. 1 in the main text but over a wider energy range. DFT+$U$ has been used to increase the semiconductor band gaps to match experiment, which is reflected in the figure. One can see that the surface bands disperse through the semiconductor band gap, and we note that there is no energy overlap between the bulk bands and the surface bands in any of the materials. In the case of Pb/SiC(0001), another isolated band lies within the bulk gap at approximately $-0.75$ eV. Orbital analysis of this band reveals it is of adatom origin, with strong Pb $p_{x}$ and $p_{y}$ character; hence, it was ignored for calculating the size of the bulk gap in Sec. II.
@@ Appendix B Electron-phonon coupling
In Tab. 7 we compare EPC calculations performed with and without DFT+$U$ for the X/Si(111) materials, in-plane phonon modes, and the effect of SOC in phonon calculations. We find that the inclusion of DFT+$U$ causes a reduction in the EPC strength $\lambda$. Notably, the deformation potential $\mathcal{D}$ is mostly unaffected, with the largest change $\Delta\mathcal{D}=0.010\,$eV/Å  in Pb/Si(111) coupled to the band 1 $K$ mode. The source of this reduction can be mostly attributed to the general increase in phonon mode frequency seen in both materials for a given mode, as $\lambda\sim 1/\omega^{2}$. For both Sn/Si(111) and Sn/SiC(0001) we test phonon modes with eigenvector corresponding to in-plane motion. For these calculations we find no discernible change in the effective band structure as compared to the unperturbed unit cell and thus minimal coupling. It is unsurprising that out-of-plane motion will have the largest effect on the band structures when considering hybridization of Wannier functions with the bulk. As discussed in the main text, the reduction in hybridization corresponds to significant changes in tight-binding hopping parameters. Changes in the adatom height above the substrate should have the largest effect on hybridization, while in-plane motion should leave this largely unchanged.
Material Mode $\lambda$ $\omega\,{\rm(meV)}$ $\mathcal{D}\,({\rm eV}/\text{\AA})$
@@
Figure 8: DFT+$U$ bandstructures for a Sn/Ge(111) and b Pb/Ge(111). A zoomed in view of the surface bands along with tight-binding models are shown for c Sn/Ge(111) and d Pb/Ge(111). Up to 10th nearest neighbor terms have been used for the tight-binding models. Adding further nearest neighbor terms gives better matches around the $\Gamma$-point. Fermi surfaces have been included as insets, with the Fermi surface of Sn/Ge(111) sitting very close to the van Hove singularity.
Appendix C Germanium substrate
Hopping Sn/Ge(111) Pb/Ge(111) SOC Sn/Ge(111) Pb/Ge(111) SOC Sn/Ge(111) Pb/Ge(111) SOC Sn/Ge(111) Pb/Ge(111)
$t_{0}$ (eV) $0.241074$ $0.122352$
$t_{1}$ (eV) $0.041324$ $0.043321$ $\alpha_{1}/t_{1}$ $0.019537$ $0.067036$ $\beta_{1}/t_{1}$ $J_{1}/t_{1}$
$t_{2}/t_{1}$ $-0.484682$ $-0.400106$ $\alpha_{2}/t_{1}$ $-0.050995$ $0.064605$ $\beta_{2}/t_{1}$ $J_{2}/t_{1}$ $0.042881$ $0.042658$
$t_{3}/t_{1}$ $0.138612$ $0.118972$ $\alpha_{3}/t_{1}$ $0.001301$ $0.000853$ $\beta_{3}/t_{1}$ $J_{3}/t_{1}$
$t_{4}/t_{1}$ $-0.094473$ $-0.073313$ $\alpha_{4}/t_{1}$ $0.009354$ $-0.010424$ $\beta_{4}/t_{1}$ $-0.000007$ $0.003841$ $J_{4}/t_{1}$ $0.008058$ $0.006787$
$t_{5}/t_{1}$ $-0.013309$ $-0.002355$ $\alpha_{5}/t_{1}$ $0.002634$ $-0.002310$ $\beta_{5}/t_{1}$ $J_{5}/t_{1}$
$t_{6}/t_{1}$ $-0.090335$ $-0.078438$ $\alpha_{6}/t_{1}$ $-0.005791$ $0.007177$ $\beta_{6}/t_{1}$ $J_{6}/t_{1}$ $0.007211$ $0.006879$
$t_{7}/t_{1}$ $-0.016891$ $-0.009418$ $\alpha_{7}/t_{1}$ $0.001802$ $-0.001762$ $\beta_{7}/t_{1}$ $0.000073$ $-0.000223$ $J_{7}/t_{1}$ $0.001162$ $0.001339$
$t_{8}/t_{1}$ $-0.018948$ $-0.015835$ $\alpha_{8}/t_{1}$ $0.001201$ $0.001412$ $\beta_{8}/t_{1}$ $J_{8}/t_{1}$
$t_{9}/t_{1}$ $-0.025869$ $-0.021675$ $\alpha_{9}/t_{1}$ $0.001628$ $-0.001869$ $\beta_{9}/t_{1}$ $-0.000016$ $0.000308$ $J_{9}/t_{1}$ $0.001960$ $0.001962$
$t_{10}/t_{1}$ $-0.013261$ $-0.010480$ $\alpha_{10}/t_{1}$ $0.000765$ $-0.000824$ $\beta_{10}/t_{1}$ $-0.000045$ $0.000072$ $J_{10}/t_{1}$ $0.000121$ $-0.000023$
Table 8: Tight-binding parameters for X/Ge(111). Up to 10th nearest neighbor terms are listed for the hopping, conventional Rashba, radial Rashba, and valley-Zeeman parameters.
For completeness we performed the same calculations in the main text to the materials Sn/Ge(111) and Pb/Ge(111). First-principles calculations for these materials can be found in the literature, see for instance Refs. Carpinelli et al., 1996, 1997; Schuwalow et al., 2010; Tresca and Calandra, 2021; Profeta and Tosatti, 2007.
The use of DFT+$U$ is essential for X/Ge(111) as germanium is predicted as metallic in conventional DFT. We use values of $U_{\rm Sn/Ge}=-7.25\,{\rm eV}$ and $U_{\rm Pb/Ge}=-7.86\,{\rm eV}$ to match germanium’s $\Gamma$-point gap.
The structures were relaxed following the method for X/SiC(0001) outlined in the main text, with $10\,{\rm\AA}$ of vacuum used for both materials. We find that the adatoms sit above the substrate at heights of $h=1.95\,{\rm\AA}$ for Sn/Ge(111) and $h=1.87\,{\rm\AA}$ for Pb/Ge(111).
The surface bands are shown in Fig. 8 and have slightly larger bandwidths than the materials in the main text with $W/t_{1}=11.10$ $(W=0.46\,\rm{eV})$ for Sn/Ge(111) and $W/t_{1}=10.58$ $(W=0.46\,\rm{eV})$ for Pb/Ge(111).
The surface bands in both X/Ge(111) material show a dip around the $\Gamma$-point. A denser $16\times 16\times 1$ $\Gamma$-centered grid was required to capture this feature during the Wannierization process. Tight-binding parameters including all SOC terms are listed in Tab. 8. The tight-binding models plotted in Fig. 8 c,d have up to 10th nearest neighbor contributions as listed in Tab. 8. This gives a good fit around the $K$ and $M$ high-symmetry points, and adding further nearest neighbor terms gives increasingly better fits around the $\Gamma$-point. Rashba SOC is weaker compared to the materials discussed in the main text, with $\alpha_{1}/t_{1}=0.02$ for Sn/Ge(111) and $\alpha_{1}/t_{1}=0.07$ for Pb/Ge(111). Both materials feature quite large $t_{2}/t_{1}$ and $\alpha_{2}/t_{1}$ values (compared to the other materials), leading to the dip around the $\Gamma$-point and another crossing of the bands along the $\Gamma K$ path. Compared to X/Si(111) and X/SiC(0001) the van Hove singularities sit much closer to the Fermi surface, especially in the case of Sn/Ge(111).
Figure 9: Phonon dispersions for a Sn/Ge(111) and b Pb/Ge(111). The band corresponding to out-of-plane motion of the adatom sits above the two in-plane bands. Bulk modes sit at lower frequencies compared to X/Si(111) and X/SiC(0001) due to the higher atomic mass of Ge.
Phonon dispersions are calculated using $2\times 2$ supercells and DFT+$U$ and are displayed in Fig. 9. In contrast to X/Si(111) and X/SiC(0001), the X/Ge(111) materials have out-of-plane surface bands sitting at higher frequencies than the in-plane modes. In addition, the more massive germanium atoms lead to the bulk modes appearing at lower frequencies.
We calculate electron phonon coupling strengths for these materials following the same method outlined in the main text, with results listed in Tab. 9. We focus on the out-of-plane phonon modes and find that the EPC strengths $\lambda$ are too low for conventional superconductivity. The density of states at the Fermi energy is comparatively much higher in the germanium materials as the van Hove singularity lies very close to the Fermi level, though this is balanced by the low deformation potentials $\mathcal{D}$.
Material Mode $\lambda$ $\omega\,{\rm(meV)}$ $\mathcal{D}\,({\rm eV}/\text{\AA})$ $N(E_{F})$
(band ##\##)
Sn/Ge(111) $K$ (3) 0.013 6.957 0.057 $5.57\frac{1}{\rm eV}$
$M$ (3) 0.046 6.873 0.106
Pb/Ge(111) $K$ (3) 0.025 5.920 0.111 $3.63\frac{1}{\rm eV}$
$M$ (3) 0.051 5.962 0.158
Table 9: Electron-phonon coupling parameters for X/Ge(111): phonon mode and band index, EPC strength $\lambda$, frequency $\omega$ of the mode, deformation potential $\mathcal{D}$ and DOS at $E_{F}$.